OpenStudy (anonymous):

An object moves along the line so that it's velocity at time "t" is v(t) = 3t^2 - 24t + 36 ft/s. Find the displacement and the total distance traveled by the object for -1< (or equal) t < (or equal) 9 An object moves along the line so that it's velocity at time "t" is v(t) = 3t^2 - 24t + 36 ft/s. Find the displacement and the total distance traveled by the object for -1< (or equal) t < (or equal) 9 @Mathematics

OpenStudy (anonymous):

1\[\le\]t\[\le\]9

OpenStudy (anonymous):

for displacement just do \[\int\limits_{1}^{9}v.dt = t^3 -12t^2 +36t \] and apply the limits from 1 to 9

OpenStudy (anonymous):

this was for displacement, now for finding the distance, we will have to integrats modulous of v.dt in the given interval

OpenStudy (anonymous):

May I see the complete work out if possible please